evaluate ∫ (3x + 4)^2 dx

evaluate
∫ (3x + 4)^2 dx

Grade:12

1 Answers

Jitender Singh IIT Delhi
9 years ago
Ans:
Hello Student,
Please find answer to your question below

I = \int (3x+4)^2dx
I = \int (9x^2+16+24x)dx
I = \int 9x^2dx+\int 16dx+\int 24xdx
I = 3x^3+\int 16dx+\int 24xdx
I = 3x^3+16x+\int 24xdx
I = 3x^3+16x+12x^2+c